OpenStudy (anonymous):

Please Help!!! find the slope and y-intercept of the graph of this graph 6y=-8x-18

OpenStudy (anonymous):

The slope would be 4\3 and the y-intercept would be -3. You can figure this out by putting the equation in y=mx+b form

OpenStudy (johnweldon1993):

@MathStudent33 is close... Like he said you can solve this equation for 'y' and you can see the slope and y-intercept...so lets do that... \[\large 6y = -8x -18\] To make this just y = something....we divide both sides of this equation by 6... \[\large \frac{6y}{6} = \frac{-8x}{6} -\frac{18}{6}\] So now we have \[\large y = -\frac{8}{6}x - 3\] Which yes can be simplified into \[\large y = -\frac{4}{3}x - 3\] So your slope is -4/3 and y-intercept is -3 *Just forgot the minus sign*

OpenStudy (anonymous):

Aww man I was close but thanks for the correction @johnweldon1993 and actually *she